OpenStudy (rosales0228):

Will give medal and fan! Solve 3x - 5y = 17 9x - 15y = -4

OpenStudy (rosales0228):

(-1, -4) (4, -1) No Solutions Infinitely Many Solutions I think its No solutions... but don't know if im right

Vocaloid (vocaloid):

well, show me how you got that answer and I'll tell you if you're right or not

OpenStudy (michele_laino):

hint: If I multiply both sides of the first equation by \(3\), I get the subsequent system: \[\huge \left\{ \begin{gathered} 9x - 15y = 51 \hfill \\ 9x - 15y = - 4 \hfill \\ \end{gathered} \right.\] As we can see the left sides, of such equations, are equal each other, nevertheless the right sides are different each from other, so, what can you conclude?